# Zeroth-order log-concave sampling

### v1

@@ 5.2.1 Proximal sampler for tilted Gaussian
Consider the annealing distribution $\mu=\mu_{\sigma^{2},\rho}$ above, given access to an evaluation oracle for $V$. For $q_{0}>1$, let $\mu_{k}$ be the law of the $k$-th iterate of $\mathsf{PS}_{\textup{ann}}$, initialized from $\mu_{0}$ toward the target $\mu$, with step size $h\asymp(d^{2}\,\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu))^{-1}$ and threshold $\tau\asymp\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)\,e^{\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)}$. Then, for any $\varepsilon>0$ and $q>1$, we can ensure $\mathcal{R}_{q}(\mu_{N}\mathbin{\|}\mu)\leq\varepsilon$ after $N\asymp d^{2}(\sigma^{2}\vee 1)\log\frac{q\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)}{(2\wedge q_{0}-1)\,\varepsilon}$ iterations, using
$2\tau N=\widetilde{O}\bigl{(}d^{2}(\sigma^{2}\vee 1)\,e^{\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)}\log\frac{q}{(2\wedge q_{0}-1)\,\varepsilon}\bigr{)}\qquad\text{evaluation queries in expectation}\,.$
Mixing analysis.
Using the Bakry–Émery criterion and bounded perturbation, one can show that $C_{\mathsf{LSI}}(\mu_{i})\lesssim\sigma_{i}^{2}\vee 1$ (see [KV25b, Lemma 3.3]). Thus, when $\mathsf{PS}_{\textup{ann}}$ starts from $\mu_{0}^{V}$, by Corollary 1.3, it needs at most $N\lesssim h^{-1}\,(\sigma^{2}\vee 1)\log\frac{q\mathcal{R}_{2\wedge q_{0}}(\mu_{0}^{V}\mathbin{\|}\mu^{V})}{(2\wedge q_{0}-1)\,\varepsilon}$ iterations to obtain a sample $X_{N}$ such that $\mathcal{R}_{q}(\operatorname{law}X_{N}\mathbin{\|}\mu^{V})\leq\varepsilon$.
Per-step query complexity.
Denoting by $\ell(w)$ the success probability of each trial at $w=(y,s)$, we recall the density of $\mu^{W}=\mu^{V}*\gamma_{h}$ from [KV25b, Lemma 3.7]:
$\mu^{W}(y,s)=\frac{r^{d/2}\ell(y,s)}{\int_{\bar{\mathcal{K}}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2}-\rho t)\,\mathrm{d}x\mathrm{d}t}\,\exp\bigl{(}-\frac{1}{2\tau\sigma^{2}}\,|y_{r}|^{2}\bigr{)}\exp\bigl{(}-\rho s+\frac{1}{2}\,\rho^{2}h\bigr{)}\,.$
Moreover, $\mu^{W}$ has a concentration property (see [KV25b, Lemma 3.8]) that if $\delta\geq 24hd$,
$\mu^{W}(\mathbb{R}^{d+1}\backslash\widetilde{\mathcal{K}})\leq\exp\bigl{(}-\frac{\delta^{2}}{2h}+24\delta d+hd^{2}\bigr{)}\quad\text{for}\quad\widetilde{\mathcal{K}}=\mathopen{}\mathclose{{}\left[\begin{array}[]{cc}r^{-1}I_{d}\\ &1\end{array}}\right]\bar{\mathcal{K}}_{\delta}+\mathopen{}\mathclose{{}\left[\begin{array}[]{c}0\\ \rho h\end{array}}\right]\,.$
Choosing $h=(24^{2}cd^{2})^{-1}$ and $\delta=(1+\nicefrac{{1}}{{24c}})\,d^{-1}$ for some $c\geq 1$, we can ensure that $\delta\geq 24hd$ and $\mu^{W}(\mathbb{R}^{d+1}\backslash\widetilde{\mathcal{K}})\leq e^{-c/2+1}$. There is also an expansion lemma (see [KV25b, Lemma 3.9]): for $\rho\in(0,d]$,
$r^{-d/2}\int_{\bar{\mathcal{K}}_{s}}e^{-\frac{1}{2r\sigma^{2}}\,|z|^{2}-\rho l}\,\mathrm{d}z\mathrm{d}l\leq 2\exp(hd^{2}+24sd)\int_{\bar{\mathcal{K}}}e^{-\frac{1}{2\sigma^{2}}\,|z|^{2}-\rho l}\,\mathrm{d}z\mathrm{d}l\,.$
Using change of variables and this expansion lemma,
$\frac{\int_{\widetilde{\mathcal{K}}}r^{d/2}\exp(-\frac{1}{2\tau\sigma^{2}}\,|y_{r}|^{2}-\rho s+\frac{1}{2}\,\rho^{2}h)}{\int_{\bar{\mathcal{K}}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2}-\rho t)}=\frac{r^{-d/2}\int_{\bar{\mathcal{K}}_{\delta}}\exp(-\frac{1}{2r\sigma^{2}}\,|y|^{2}-\rho s-\frac{1}{2}\,\rho^{2}h)}{\int_{\bar{\mathcal{K}}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2}-\rho t)}\leq e^{27}$ (5.3)
Similarly for earlier analyses, the failure probability of one stage in the $k$-th iteration can be bounded by $1/2$ under suitable choices of $h$ and $\tau$. For a conjugate pair $(p,q)$ with $p,q>1$, we have $\mathbb{E}_{\mu_{k}^{W}}[(1-\ell)^{\tau}]\leq M_{q}\,\{\mathbb{E}_{\mu_{k}^{W}}[(1-\ell)^{\tau p}]\}^{1/p}$ and
$\mathbb{E}_{\mu_{k}^{W}}[(1-\ell)^{\tau p}]=\int_{\mathbb{R}^{d+1}\backslash\widetilde{\mathcal{K}}}\cdot+\int_{\widetilde{\mathcal{K}}\cap[\ell\tau\geq\log(3M_{q})]}\cdot+\int_{\widetilde{\mathcal{K}}\cap[\ell\tau<\log(3M_{q})]}=:\mathsf{A}+\mathsf{B}+\mathsf{C}\,.$
Under $c\asymp\log(3M_{q})^{p}$ and $\tau\asymp M_{q}^{p}\log(3M_{q})$, the RHS can be bounded by $(1/2M_{q})^{p}$ as seen in
$\mathsf{A}$ $\leq\mu^{W}(\mathbb{R}^{d+1}\backslash\widetilde{\mathcal{K}})\leq e^{-c/2+1}\,,$
$\mathsf{B}$ $\leq\int_{\widetilde{\mathcal{K}}\cap[\ell\tau\geq\log(3M_{q})]}e^{-\ell\tau p}\,\mathrm{d}\mu^{W}\leq\bigl{(}\frac{1}{3M_{q}}\bigr{)}^{p}\,,$
$\mathsf{C}$ $\leq\int_{\widetilde{\mathcal{K}}\cap[\ell\tau<\log(3M_{q})]}\frac{r^{d/2}\ell(y,s)\exp(-\frac{1}{2r\sigma^{2}}\,|y_{r}|^{2})\exp(-\rho s+\frac{1}{2}\,\rho^{2}h)}{\int_{\bar{\mathcal{K}}}\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2}-\rho t)}$
[…]
Thus, $\mathbb{E}_{\mu_{k}^{W}}[(1-\ell)^{\tau}]\leq 1/2$, and per-iteration complexity is bounded by $2\tau$. Combining the mixing analysis and per-iteration analysis proves the query complexity of $\mathsf{PS}_{\textup{ann}}$ (Lemma 5.1).
5.2.2 Query complexity of Rényi-warmness generation
We first recall a result on closeness of two log-concave distributions from [KV25a, Lemma 4.4].
Lemma 5.2.
Let $\mathrm{d}\nu\propto e^{-V}\,\mathrm{d}x$ and $\mathrm{d}\mu\propto e^{-(1+\alpha)\,V}\,\mathrm{d}x$ be log-concave probability measures over $\mathbb{R}^{d}$. For $q>1$ and $\delta>0$ such that $01-q\delta>$ 0 ,
@@ 5.2.2 Query complexity of Rényi-warmness generation
At initialization, $\mu_{0}$ can be sampled by rejection sampling with $O(1)$ queries in expectation (see [KV25b, Lemma 5.6]). In Phase I, any consecutive distributions are $O(1)$-close in $\mathcal{R}_{2}$ by Lemma 5.2. Since there are $O(d^{1/2}\log d)$ many phases, with each one requiring $\widetilde{O}(d^{2})$ queries, the query complexity of Phase I is $\widetilde{O}(d^{2.5})$. Meanwhile, we relay $\mathcal{R}_{2}$-warmness through the weak triangle inequality (as done in the proof of Theorem 1.5).
@@
In Phase II, closeness during the simultaneous annealing (called the outer annealing hereafter) is justified again by Lemma 5.2, and $\mathsf{PS}_{\textup{ann}}$ requires $\widetilde{O}(d^{2})$ queries for each outer annealing. Each outer annealing is followed by $O(D/d^{1/2})$ many inner phases, where $\sigma^{2}$ goes back to $1$ again. Closeness within the inner annealing follows from Lemma 4.1, and $\mathsf{PS}_{\textup{ann}}$ requires $\widetilde{O}(d^{2})$ queries. Thus, completion of one outer annealing (combined with the inner annealing) requires $\widetilde{O}(d^{2}+d^{1.5}D)$ queries. As there are $O(d^{1/2}\log d)$ outer phases, the total complexity of Phase II is $\widetilde{O}(d^{2.5}+d^{2}D)$.
@@
In Phase III, closeness of annealing distributions follow again from Lemma 4.1. Since doubling of initial $\sigma^{2}$ requires $O(D/\sigma)$ phases, with phase requiring $O(d^{2}\sigma^{2})$ queries from an $O(1)$-warm start, each doubling uses $\widetilde{O}(d^{2}\sigma D)$ queries. As $\sigma^{2}$ is increased until $\sigma^{2}\lesssim qD\,(\Lambda^{1/2}\vee 1)\log^{1/2}d$, it takes Phase III $\widetilde{O}(q^{1/2}d^{2}D^{3/2}(\Lambda^{1/4}\vee 1))$ queries to obtain a sample whose law is $O(1)$-close to $\mu_{\text{last}}$ in $\mathcal{R}_{2}$-divergence. Due to hypercontractivity of $\mathsf{PS}$ again, iterating $\mathsf{PS}_{\textup{ann}}$ toward $\mu_{\text{last}}$ $O(d^{2}\sigma_{\text{last}}^{2}\log_{2}q)$ times returns a sample $Z^{*}$ such that $\mathcal{R}_{2q}(\operatorname{law}Z^{*}\mathbin{\|}\mu_{\text{last}})=O(1)$. Hence, Phase III uses $\widetilde{O}(qd^{2}D^{3/2}\,(\Lambda^{1/4}\vee 1))$ queries in total. Also, by the early-stopping result and weak triangle inequality, we have $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\bar{\pi})=O(1)$. Since $\mathcal{R}_{\infty}(\bar{\pi}\mathbin{\|}\pi)=O(1)$, we obtain that $\mathcal{R}_{q}(\operatorname{law}Z^{*}\mathbin{\|}\pi)=O(1)$.

### v2

@@ 5.2.2 Query complexity of Rényi-warmness generation
Let $\mathrm{d}\nu\propto e^{-V}\,\mathrm{d}x$ and $\mathrm{d}\mu\propto e^{-(1+\alpha)\,V}\,\mathrm{d}x$ be logconcave probability measures over $\mathbb{R}^{d}$. For $q>1$ and $\delta>0$ such that $1-q\delta>0$,
@@
We set $\eta_{\exp}=\varepsilon_{\exp}=\varepsilon<1/100$ and consider $q_{0}$ and $q\geq 2\vee\widetilde{O}(\log\frac{dD}{\varepsilon})$ (to satisfy (5.2)). At initialization, $\mu_{0}$ can be sampled by rejection sampling with $O(1)$ queries in expectation (see [KV25b, Lemma 5.6]). In Phase I, any consecutive distributions satisfy $\mathcal{R}_{2q_{0}}(\mu_{i}\mathbin{\|}\mu_{i+1})\leq 1$ by Lemma 5.3. Since there are $O(q_{0}^{1/2}d^{1/2}\log d)$ many phases, with each one requiring $\widetilde{O}(d^{2}\log^{2}q_{0}m)$ queries from an $O(1)$-warm start, the query complexity of Phase I is $\widetilde{O}(q_{0}^{1/2}d^{2.5}\log^{2}q_{0}m)$. Meanwhile, we relay $\mathcal{R}_{q_{0}}$-warmness through the weak triangle inequality (as done in the proof of Theorem 1.5).
@@
In Phase II, closeness during the simultaneous annealing (called the outer annealing hereafter) is justified again by Lemma 5.3, and $\mathsf{PS}_{\textup{ann}}$ requires $\widetilde{O}(d^{2}\log^{2}q_{0}m)$ queries for each outer annealing. Each outer annealing is followed by $O(q_{0}^{1/2}D/d^{1/2})$ many inner phases, where $\sigma^{2}$ goes back to $1$ again. Closeness within the inner annealing follows from Lemma 4.2, and $\mathsf{PS}_{\textup{ann}}$ requires $\widetilde{O}(d^{2}\log^{2}q_{0}m)$ queries. Thus, completion of one outer annealing (combined with the inner annealing) requires $\widetilde{O}((d^{2}+q_{0}^{1/2}d^{1.5}D)\log^{2}q_{0}m)$ queries. As there are $O(q_{0}^{1/2}d^{1/2}\log d)$ outer phases, the total complexity of Phase II is $\widetilde{O}((q_{0}^{1/2}d^{2.5}+q_{0}d^{2}D)\log^{2}q_{0}m)$.
@@
In Phase III, closeness of annealing distributions follow again from Lemma 4.2. Since doubling of initial $\sigma^{2}$ requires $O(q_{0}D/\sigma)$ phases, with phase requiring $\widetilde{O}(d^{2}\sigma^{2}\log^{2}q_{0}m)$ queries from an $O(1)$-warm start, each doubling uses $\widetilde{O}(q_{0}d^{2}\sigma D\log^{2}q_{0}m)$ queries. As $\sigma^{2}$ is increased while $\sigma^{2}\lesssim qD\,(\Lambda^{1/2}\vee 1)\log^{1/2}d$, it takes Phase III $\widetilde{O}(q^{1/2}q_{0}d^{2}D^{3/2}\,(\Lambda^{1/4}\vee 1)\log^{2}q_{0}m)$ queries to obtain a sample whose law is $O(1)$-close to $\mu_{m}$ in $\mathcal{R}_{q_{0}}$-divergence. Due to hypercontractivity of $\mathsf{PS}$ again, iterating $\mathsf{PS}_{\textup{ann}}$ toward $\mu_{m}$ $O(d^{2}\sigma_{m}^{2}\log^{2}qm)$ times returns a sample $Z^{*}$ such that $\mathcal{R}_{2q}(\operatorname{law}Z^{*}\mathbin{\|}\mu_{m})\leq 1$. Hence, Phase III uses
$\widetilde{O}\bigl(q^{1/2}q_{0}d^{2}D^{3/2}\,(\Lambda^{1/4}\vee 1)\log^{2}q_{0}m+qd^{2}D\,(\Lambda^{1/2}\vee 1)\log^{2}qm\bigr)=\widetilde{O}\bigl(qd^{2}D^{3/2}(\Lambda^{1/4}\vee 1)\log^{2}qm\bigr)$
queries in total. Also, by the early-stopping result and weak triangle inequality, we have $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\bar{\pi})\leq\frac{1.1}{10}+\frac{1}{2}$. Since $\mathcal{R}_{\infty}(\bar{\pi}\mathbin{\|}\pi)\leq\log 2$, we obtain that $\mathcal{R}_{q}(\operatorname{law}Z^{*}\mathbin{\|}\pi)\leq 1$.